rose invested part of her money at 8% per year and twice of that amount at 5% per year. if the returns last year from the two investments were 7200 how much did rose invest at each rate?
Answer by mananth(16949)

-------------- 8.00% per annum ------------- x
------------- 5.00% per annum ------------y y=2x
total investment 50000
Interest----- $7,200.00

8.00% x + 5.00% * 2 x = $7,200.00
0.08 x + 0.05 * 2 x = $7,200.00
0.08 x + 0.1 x = $7,200.00
0.18 x = $7,200.00
/ 0.18
x= $40,000.00 @ 7%
2x = $80,000.00 @5%


CHECK
$3,200.00 + $4,000.00 = $7,200.00